Grade 10Discuss with Askiitians Tutors

There are 1000 people standing in a circle. they are numbered 1 to 1000. 2 is on the left of 1,3 is on the left of 2 and so on. Starting from 1, each person kills the person on the left. Who will be the last person standing.

Rupam Mandal
9 Years agoGrade 10

1 Answer

Arun
6 Years ago
I tried in this way using the formula of a guest
 
logarithm of 1000 base 2 = 9.9~10
 
2^10 = 1024
 
now of soldiers = 1000
 
1000-1024=-24
 
solider at 24th position in anti clock wise direction survives
 
that is 1000-24+1=977
 
I tried in this way using the formula of a guest
 
logarithm of 1000 base 2 = 9.9~10
 
2^10 = 1024
 
now of soldiers = 1000
 
1000-1024=-24
 
solider at 24th position in anti clock wise direction survives
 
that is 1000-24+1=977